OpenStudy (anonymous):

Evaluate A 2-B 2 for A = 3 and B = -4.

OpenStudy (akashdeepdeb):

I am guessing you have been asked to find this: \[A^2 - B^2~~~~~~~~~~ When~~~ A=3, ~~~~B=-4\] So, here , all you need to do is plug in the values of A and B given to you! :D \[(3)^2 - (-4)^2 = 9 - 16\] Understood?

OpenStudy (anonymous):

is the answer -2

OpenStudy (akashdeepdeb):

9-16 is not -2 is it?

OpenStudy (anonymous):

mybad -7 im sorry

OpenStudy (akashdeepdeb):

-7 is correct. :)